OpenStudy (anonymous):

Two functions, C and D, are described as follows: Function A y = 9x + 4 Function B The slope is 3 and the y-intercept is 4. How much more is the slope of function A than the slope of function B? 2 3 6 9

OpenStudy (anonymous):

please anyone help me

OpenStudy (anonymous):

@narissa

OpenStudy (anonymous):

@159507nb

OpenStudy (anonymous):

@MollyhatesMath

OpenStudy (anonymous):

@Thefaceless

OpenStudy (anonymous):

please helpguys please

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

y=mx+b FunctionA y=9x+4 m=9, b=4

OpenStudy (anonymous):

9-3=? thats your answer